Muhammad Raza company is a producer of pastries. The company hires an economist to determine the demand for its product. After months of hard work, the analyst tells the company that demand for the fim's pastries(Qx) is given by the following
equations:
Qx = 20000 - 10000 Px + 10 I + 1000 Pc
Where Px is the price charged for Raza pastries, I is income per capita and Pc is the price of books from competing publishers. Using this information, the company's managers want to:
(a) Determine what effect a price increase would have on total revenues.
(b) Evaluate how sale of pastries would change during a period of raising income.
(c) Assess the probable impact if competing producers raise their prices. Assume that the initial values of Px = $8, I = $18000 and Pc = S10
1
Expert's answer
2021-07-12T15:54:46-0400

Muhammad Raza Company is producer of pastries.

Qx = 20000-10000Px+10I+1000Pc


Qx = 20000-10000(8)+10(18000)+1000(10)


Qx = 20000-80000+180000+10000= $130000



Where Px is the price charged for Raza pastries, I is the income per capita and Pc is the price of pastries form competing producer.

(i) The effect a price increase would have on total revenues will be negative, because the demand is elastic and total revenue will decrease.

(ii) Sale of pastries would decrease during a period of raising income, because the competing good is more preferable for the customers.

(iii) If competing producers raise their prices, the demand for this product will increase.